# TokLIP: Marry Visual Tokens to CLIP for Multimodal Comprehension and Generation

### v1

@@ 1 Introduction
A unified autoregressive model with end-to-end next multimodal token prediction is considered the way toward the ChatGPT moment for the multimodal world. Pioneering works, such as Chameleon [50] and Emu3 [58], train a single Transformer with vector-quantized (VQ) visual tokens and word tokens on multimodal sequences. Though scalable, the models suffer from expensive computational overhead for training to converge and weak comprehension performance.
VILA-U et al. [61, 42, 71] highlight that the issue lies in the fact that VQ tokens lack high-level semantics, which CLIP [43] possesses and has proven effective for LLaVA [34, 35]-like frameworks designed solely for comprehension purposes. They attempt to solve the problem via discretizing high-level semantical features, that is, training a continuous-to-discrete visual tokenizer with both reconstruction and text alignment objectives, illustrated in Figure 1 (a). However, the conflict between two distinct objectives and the information loss of visual semantics caused by quantization operation pose new challenges for properly unifying multimodal comprehension and generation.
@@
Specifically, the TokLIP tokenizer consists of an off-the-shelf low-level VQ tokenizer, such as VQGAN [14, 47], and a ViT-based token encoder with causal attention. An image is first discretized at a low level, after which the visual code embeddings are fed into the token encoder to capture high-level continuous semantics. The token encoder is pre-trained with only semantic supervision, specifically through text alignment and CLIP distillation, and is subsequently equipped with the autoregressive Transformer for end-to-end multimodal token prediction.
@@
We empirically validate the effectiveness of TokLIP by comparing it with other tokenizers and integrating it with large language models (LLMs) for autoregressive multimodal comprehension and generation. Our analysis reveals three key findings: TokLIP demonstrates exceptional image representation capabilities, utilizing less than 20% of the pretraining data required by VILA-U, while significantly outperforming established tokenizers such as VILA-U and QLIP in zero-shot ImageNet classification. Additionally, TokLIP retains the original reconstruction capabilities of the low-level VQ tokenizer without necessitating specific optimization for generation and quantization. (2) Multimodal comprehension at significantly reduced training cost: When integrated with LLMs for comprehension tasks, TokLIP achieves comparable results using only 3% of the training data required by SynerGen-VL. This reveals that TokLIP unleashes the potential of discrete token-based multimodal LLMs to efficiently deliver advanced comprehension performance similar to that of the LLaVA series. (3) Synergistic autoregressive image generation: We demonstrate that high-level semantic features complement low-level code embeddings to enhance image generation, underscoring TokLIP’s ability to integrate diverse abstraction levels for robust multi-modal outputs.
@@ 2 Related Works
Vector quantization (VQ) is effective in processing different modalities, especially images [55, 14, 65, 25, 4, 57, 44, 52], compressing high-dimensional raw data into discrete token sequences for efficient learning. The pioneering work VQ-VAE [55] proposes to discretize continuous representations by assigning them to the closest vectors in a learnable codebook. VQGAN [14] incorporates adversarial and perceptual losses to further improve generation quality. Recent studies, such as LlamaGen [47], MAGVIT2 [38, 66] and VAR [51], demonstrate the great potential of VQ tokenizers in autoregressive image generation. Such tokenizers are naturally suited for generation tasks due to capturing low-level patterns, however, they struggle to provide necessary semantics for multimodal comprehension tasks, as evidenced in [59, 62]. In this work, we present TokLIP, which semanticizes VQ tokens to enhance multimodal comprehension while maintaining end-to-end multimodal autoregressive learning on plain VQ tokens when integrated into large multimodal models.
@@
Exploring unified modeling of multimodal comprehension and generation in the framework of large language models has gained significant attention. Despite various attempts, the core challenge remains addressing the essential conflicts of comprehension and generation tasks. There are four categories of approaches to meet the challenge: (1) External diffusion: [19, 48, 53, 20, 49, 60] utilize external pretrained diffusion models linked through semantic features for image generation, resulting in independent processing of comprehension and generation. (2) Internal diffusion: [62, 72, 46] facilitate visual-language comprehension and generation within a single model; however, their use of denoising steps during image generation tasks prevents true fusion of images and text. (3) Dual encoders: [59, 39, 7] employ a continuous semantic encoder for comprehension and a discrete encoder for generation, which incurs increased computational overhead and results in independent processing of these tasks. (4) Unified tokenizer: [33, 50, 58] pioneered this line of research by training a single autoregressive Transformer with discrete image VQ tokens and text tokens. [42, 61, 71] further improve the multimodal comprehension performance by proposing new discrete visual tokenizers with high-level semantics.
It is evident that only the "unified tokenizer" category can effectively integrate comprehension and generation tasks, along with text and image tokens, into a single autoregressive model. However, state-of-the-art methods in this area face challenges due to conflicting training objectives—namely, reconstruction losses for generation and text-alignment losses for comprehension—as well as information loss from quantization training. This complexity makes stable training particularly difficult. We contend that it is more proper to semanticize low-level discrete visual tokens, as our proposed TokLIP, rather than to discretize high-level continuous visual features, as done in previous works.
@@ 3.1 TokLIP Tokenizer
Pioneering works [50, 58] in multimodal unification adopt VQ-based low-level tokenizers to encode images for both comprehension and generation tasks. These fidelity-oriented tokenizers primarily capture basic image features, such as contours and patterns, making them ideal for generation tasks. However, visual understanding tasks require high-level semantic representations to enhance interpretability and reasoning capabilities, as evidenced in [34, 73, 2]. The lack of such semantic knowledge in VQ-based tokenizers limits their effectiveness in multimodal understanding.
@@
To solve the problem, we explore the possibility of enriching VQ-based tokenizers [14, 47] with high-level semantic knowledge. CLIP [43, 70], trained on hundreds of millions of image-text pairs, has demonstrated remarkable success in visual understanding. To this end, we propose TokLIP, a framework that integrates fidelity encoders with CLIP vision encoders to simultaneously preserve both generation and comprehension capabilities in a single tokenizer, as illustrated in Figure 2 (a). Specifically, given an input image $I\in\mathbb{R}^{H\times W\times 3}$, we first quantize it using a VQGAN encoder $\mathcal{E}_{\text{vq}}$ to obtain the code embeddings of discrete tokens $x_{\text{vq}}=\mathcal{E}_{\text{vq}}(I)\in\mathbb{R}^{(\frac{H.W}{p.p})\times\hat{d}}$, where $p$ is the downsample ratio and $\hat{d}$ is the dimension of VQGAN codebook embeddings. The code embeddings $x_{\text{vq}}$ are further mapped to CLIP input feature dimension $d$ via a multi-layer perception (MLP) layer before being fed into a token encoder $\mathcal{E}_{\text{tok}}$ which shares the same architecture as CLIP’s ViT encoder but uses causal attention. To conclude, the TokLIP tokenizer encodes the input image into semantic features $x_{\text{toklip}}\in\mathbb{R}^{(\frac{H.W}{p.p})\times d}$ from discrete to continuous, denoted as
@@
In the original CLIP architecture [43], the vision Transformer encoder employs bidirectional attention to capture the rich semantic relationships among image patches and facilitate a holistic understanding of the input. However, in the context of multimodal autoregressive learning, the bidirectional attention mechanism hinders the training paradigm of end-to-end next token prediction, where each token can only be predicted by attending to its preceding tokens. To address this discrepancy, we adopt causal attention for our token encoder, which enables end-to-end next multimodal token prediction when further equipping TokLIP with LLMs as shown in Figure 2 (b). In such a way, image and text signals, as well as comprehension and generation tasks, can be truly fused in a single causal Transformer.
@@
We emphasize that TokLIP is not a fixed model architecture but a continuously improvable framework that can be enhanced with advanced VQ-based tokenizers and powerful CLIP encoders. For instance, we can replace VQGAN with superior tokenizers such as MAGVIT-v2 [66], Cosmos [1] or large-scale pre-trained tokenizers as utilized in Emu3 [58]. Additionally, we can substitute CLIP encoders with more effective ViT-based semantic encoders like SigLIP [70], AIM [13], or DFN [15]. Notably, TokLIP can be further extended by combining it with a diffusion head, following the success of GPT-4o [6, 40], to generate higher-quality and high-resolution images.
@@ 3.3 Discussion
Another line of work aims to train a semantic tokenizer by quantizing high-level image features from the CLIP encoder into discrete tokens [61, 71]. These methods typically rely on joint training with both reconstruction and semantic losses. The key distinction between these approaches and TokLIP is that the former adopts a pipeline that transitions from continuous to discrete representations (i.e., discretize high-level CLIP features), while TokLIP follows a discrete-to-continuous way (i.e., semanticize low-level VQ tokens).
@@
To better understand the advantages of disentangled training objectives, we conduct a feature analysis for comparison, as shown in Figure 3, demonstrating that TokLIP learns better semantics than VILA-U. We primarily compare the feature similarity between tokenizers and their initialized SigLIP [70, 54] encoders. Specifically, we extract the normalized features $x_{\text{toklip}}$/$x_{\text{vila-u}}$ and $x_{\text{siglip}}$ for a single input image and compute the cosine similarity between $x_{\text{toklip}}$/$x_{\text{vila-u}}$ and $x_{\text{siglip}}$. Our results show that TokLIP achieves a similarity score of 0.83, outperforming VILA-U, which scores 0.75. We further visualize the normalized features using t-SNE in Figure 3 (a). On the left, the VILA-U features are far from the SigLIP-ViT-L/16 features, indicating that the feature spaces differ significantly. In contrast, TokLIP features are much closer to SigLIP2-So/16. These results suggest that the training process in VILA-U drastically alters the feature representations compared to the semantic encoder. We attribute this to the conflict between the reconstruction and semantic objectives and the information loss during the quantization process.
Additionally, we analyze the image representations used for retrieval and classification tasks, which are achieved by pooling image features $x_{\text{siglip}}$ through an attention pooling head in SigLIP [70]. We randomly select 1,500 images from the MSCOCO test set and visualize these representations using t-SNE in Figure 3 (b). Notably, the representations from VILA-U are relatively distant from SigLIP, while TokLIP shares nearly the same distribution as SigLIP2. This aligns with TokLIP’s stronger retrieval performance demonstrated in Section 4.2. These analyses further show that our pipeline learns more similar semantics to pre-trained CLIP ViTs, thus achieving superior comprehension performance.
@@
Pretrained on large-scale image-text pairs, CLIP [43, 54, 31] has achieved remarkable success in capturing high-level semantic features of images, making it highly effective for feature extraction in multimodal large language models (LLMs). Numerous studies [34, 73, 2, 37, 32] leverage CLIP’s continuous features by combining them with text tokens in LLMs and training on vision-language data to enhance performance in comprehension tasks. However, while CLIP excels at encoding high-level semantics, it often overlooks low-level structures and textural details, which are crucial for visual generation tasks. Models such as SEED-X [20], DreamLLM [12], and MetaMorph [53] incorporate CLIP’s continuous features with text tokens for autoregressive prediction but still rely on diffusion models [41] for generation inference to compensate for the lack of low-level details. In contrast, TokLIP proposes a unified tokenizer that integrates both low-level features and high-level semantics, enabling direct tokenization with CLIP encoders and allowing for more effective handling of visual generation tasks with VQGAN decoder.

### v2

@@ 1 Introduction
A unified autoregressive model with end-to-end next multimodal token prediction is considered the way toward the ChatGPT moment for the multimodal world. Pioneering works, such as Chameleon (Team, 2024) and Emu3 (Wang et al., 2024b), train a single Transformer with vector-quantized (VQ) visual tokens and word tokens on multimodal sequences. Though scalable, the models suffer from expensive computational overhead for training to converge and weak comprehension performance.
VILA-U et al. (Wu et al., 2024b; Qu et al., 2024; Zhao et al., 2025) highlight that the issue lies in the fact that VQ tokens lack high-level semantics, which CLIP (Radford et al., 2021) possesses and has proven effective for LLaVA (Liu et al., 2023, 2024b)-like frameworks designed solely for comprehension purposes. They attempt to solve the problem via discretizing high-level semantical features, that is, training a continuous-to-discrete visual tokenizer with both reconstruction and text alignment objectives, illustrated in Figure 1 (a). However, the conflict between two distinct objectives and the information loss of visual semantics caused by quantization operation pose new challenges for properly unifying multimodal comprehension and generation.
@@
Specifically, the TokLIP tokenizer consists of an off-the-shelf low-level VQ tokenizer, such as VQGAN (Esser et al., 2021; Sun et al., 2024a), and a ViT-based token encoder with causal attention. An image is first discretized at a low level, after which the visual code embeddings are fed into the token encoder to capture high-level continuous semantics. The token encoder is pre-trained with only semantic supervision, specifically through text alignment and CLIP distillation, and is subsequently equipped with the autoregressive Transformer for end-to-end multimodal token prediction.
@@
We empirically validate the effectiveness of TokLIP by comparing it with other tokenizers and integrating it with large language models (LLMs) for autoregressive multimodal comprehension and generation. Our analysis reveals three key findings: TokLIP demonstrates exceptional image representation capabilities, utilizing less than 20% of the pretraining data required by VILA-U, while significantly outperforming established tokenizers such as VILA-U and QLIP in zero-shot ImageNet classification. Additionally, TokLIP retains the original reconstruction capabilities of the low-level VQ tokenizer without necessitating specific optimization for generation and quantization. (2) Multimodal comprehension at significantly reduced training cost: When integrated with LLMs for comprehension tasks, TokLIP achieves comparable results using only 5% of the training data required by SynerGen-VL. This reveals that TokLIP unleashes the potential of discrete token-based multimodal LLMs to efficiently deliver advanced comprehension performance similar to that of the LLaVA series. (3) Synergistic autoregressive image generation: We demonstrate that high-level semantic features complement low-level code embeddings to enhance image generation, underscoring TokLIP’s ability to integrate diverse abstraction levels for robust multi-modal outputs.
@@ 2 Related Works
Vector quantization (VQ) is effective in processing different modalities, especially images (Van Den Oord et al., 2017; Esser et al., 2021; Yu et al., 2021; Lee et al., 2022; Chang et al., 2022; Wang et al., 2025a; Razavi et al., 2019; Tian et al., 2025), compressing high-dimensional raw data into discrete token sequences for efficient learning. The pioneering work VQ-VAE (Van Den Oord et al., 2017) proposes to discretize continuous representations by assigning them to the closest vectors in a learnable codebook. VQGAN (Esser et al., 2021) incorporates adversarial and perceptual losses to further improve generation quality. Recent studies, such as LlamaGen (Sun et al., 2024a), MAGVIT2 (Luo et al., 2024; Yu et al., 2023a) and VAR (Tian et al., 2024), demonstrate the great potential of VQ tokenizers in autoregressive image generation. Such tokenizers are naturally suited for generation tasks due to capturing low-level patterns, however, they struggle to provide necessary semantics for multimodal comprehension tasks, as evidenced in (Wu et al., 2024a; Xie et al., 2024a). In this work, we present TokLIP, which semanticizes VQ tokens to enhance multimodal comprehension while maintaining end-to-end multimodal autoregressive learning on plain VQ tokens when integrated into large multimodal models.
@@
Exploring unified modeling of multimodal comprehension and generation in the framework of large language models has gained significant attention. Despite various attempts, the core challenge remains addressing the essential conflicts of comprehension and generation tasks. There are four categories of approaches to meet the challenge: (1) External diffusion: (Ge et al., 2023; Sun et al., 2023; Tong et al., 2024; Ge et al., 2024; Sun et al., 2024b; Wu et al., 2023) utilize external pretrained diffusion models linked through semantic features for image generation, resulting in independent processing of comprehension and generation. Recently, (Chen et al., 2025a; Pan et al., 2025; AI et al., 2025) explore how to extend pre-trained MLLMs with DiT blocks for generation. (2) Internal diffusion: (Xie et al., 2024a; Zhou et al., 2024a; Shi et al., 2024b; Xie et al., 2025) facilitate visual-language comprehension and generation within a single model; however, their use of denoising steps during image generation tasks prevents true fusion of images and text. (3) Dual encoders: (Wu et al., 2024a; Ma et al., 2024; Chen et al., 2025c) employ a continuous semantic encoder for comprehension and a discrete encoder for generation, while (Deng et al., 2025) adopts VAE for generation and explores the scale-up of the Mixture-of-depth architecture. These works will incur increased computational overhead and result in independent processing of these tasks. (4) Unified tokenizer: (Liu et al., 2024a; Team, 2024; Wang et al., 2024b) pioneered this line of research by training a single autoregressive Transformer with discrete image VQ tokens and text tokens. (Wu et al., 2024b; Zhao et al., 2025; Han et al., 2025) further improve the multimodal comprehension performance by proposing new discrete visual tokenizers with high-level semantics. (Qu et al., 2024; Ma et al., 2025a) utilize multiple codebooks to alleviate conflicts between comprehension and generation.
It is evident that only the "unified tokenizer" category can effectively integrate comprehension and generation tasks, along with text and image tokens, into a single autoregressive model. However, state-of-the-art methods (Wu et al., 2024b; Zhao et al., 2025; Ma et al., 2025a; Qu et al., 2024) in this area face challenges due to conflicting training objectives—namely, reconstruction losses for generation and text-alignment losses for comprehension—as well as information loss from quantization training. This complexity makes stable training particularly difficult. We contend that it is more proper to semanticize low-level discrete visual tokens, as our proposed TokLIP, rather than to discretize high-level continuous visual features, as done in previous research (Wu et al., 2024b; Qu et al., 2024).
@@ 3.1 TokLIP Tokenizer
Pioneering works (Team, 2024; Wang et al., 2024b) in multimodal unification adopt VQ-based low-level tokenizers to encode images for both comprehension and generation tasks. These fidelity-oriented tokenizers primarily capture basic image features, such as contours and patterns, making them ideal for generation tasks. However, visual understanding tasks require high-level semantic representations to enhance interpretability and reasoning capabilities, as evidenced in (Liu et al., 2023; Zhu et al., 2023; Bai et al., 2023). The lack of such semantic knowledge in VQ-based tokenizers limits their effectiveness in multimodal understanding.
@@
To solve the problem, we explore the possibility of enriching VQ-based tokenizers (Esser et al., 2021; Sun et al., 2024a) with high-level semantic knowledge. CLIP (Radford et al., 2021; Zhai et al., 2023), trained on hundreds of millions of image-text pairs, has demonstrated remarkable success in visual understanding. To this end, we propose TokLIP, a framework that integrates fidelity encoders with CLIP vision encoders to simultaneously preserve both generation and comprehension capabilities in a single tokenizer, as illustrated in Figure 2 (a). Specifically, given an input image $I\in\mathbb{R}^{H\times W\times 3}$, we first quantize it using a VQGAN encoder $\mathcal{E}_{\text{vq}}$ to obtain the code embeddings of discrete tokens $x_{\text{vq}}=\mathcal{E}_{\text{vq}}(I)\in\mathbb{R}^{(\frac{H.W}{p.p})\times\hat{d}}$, where $p$ is the downsample ratio and $\hat{d}$ is the dimension of VQGAN codebook embeddings. The code embeddings $x_{\text{vq}}$ are further mapped to CLIP input feature dimension $d$ via a multi-layer perception (MLP) layer before being fed into a token encoder $\mathcal{E}_{\text{tok}}$ which shares the same architecture as CLIP’s ViT encoder but uses causal attention. To conclude, the TokLIP tokenizer encodes the input image into semantic features $x_{\text{toklip}}\in\mathbb{R}^{(\frac{H.W}{p.p})\times d}$ from discrete to continuous, denoted as
@@
In the original CLIP architecture (Radford et al., 2021), the vision Transformer encoder employs bidirectional attention to capture the rich semantic relationships among image patches and facilitate a holistic understanding of the input. However, in the context of multimodal autoregressive learning, the bidirectional attention mechanism hinders the training paradigm of end-to-end next token prediction, where each token can only be predicted by attending to its preceding tokens. To address this discrepancy, we adopt causal attention for our token encoder, which enables end-to-end next multimodal token prediction when further equipping TokLIP with LLMs as shown in Figure 2 (b). In such a way, image and text signals, as well as comprehension and generation tasks, can be truly fused in a single causal Transformer.
@@
We emphasize that TokLIP is not a fixed model architecture but a continuously improvable framework that can be enhanced with advanced VQ-based tokenizers and powerful CLIP encoders. For instance, we can replace VQGAN with superior tokenizers such as MAGVIT-v2 (Yu et al., 2023a), Cosmos (Agarwal et al., 2025) or large-scale pre-trained tokenizers as utilized in Emu3 (Wang et al., 2024b). Additionally, we can substitute CLIP encoders with more effective ViT-based semantic encoders like SigLIP (Zhai et al., 2023), AIM (El-Nouby et al., 2024), or DFN (Fang et al., 2023). Notably, TokLIP can be further extended by combining it with a diffusion head, following the success of GPT-4o (Chen et al., 2025b; OpenAI, 2025), to generate higher-quality and high-resolution images.
@@ 3.3 Discussion
Another line of work aims to train a semantic tokenizer by quantizing high-level image features from the CLIP encoder into discrete tokens (Wu et al., 2024b; Zhao et al., 2025). These methods typically rely on joint training with both reconstruction and semantic losses. The key distinction between these approaches and TokLIP is that the former adopts a pipeline that transitions from continuous to discrete representations (i.e., discretize high-level CLIP features), while TokLIP follows a discrete-to-continuous way (i.e., semanticize low-level VQ tokens).
@@
To better understand the advantages of disentangled training objectives, we conduct a feature analysis for comparison, as shown in Figure 3, demonstrating that TokLIP learns better semantics than VILA-U. We primarily compare the feature similarity between tokenizers and their initialized SigLIP (Zhai et al., 2023; Tschannen et al., 2025) encoders. Specifically, we extract the normalized features $x_{\text{toklip}}$$x_{\text{vila-u}}$ and $x_{\text{siglip}}$ for a single input image and compute the cosine similarity between $x_{\text{toklip}}$$x_{\text{vila-u}}$ and $x_{\text{siglip}}$. Our results show that TokLIP achieves a similarity score of 0.83, outperforming VILA-U, which scores 0.75. We further visualize the normalized features using t-SNE in Figure 3 (a). On the left, the VILA-U features are far from the SigLIP-ViT-L/16 features, indicating that the feature spaces differ significantly. In contrast, TokLIP features are much closer to SigLIP2-So/16. These results suggest that the training process in VILA-U drastically alters the feature representations compared to the semantic encoder. We attribute this to the conflict between the reconstruction and semantic objectives and the information loss during the quantization process.
Additionally, we analyze the image representations used for retrieval and classification tasks, which are achieved by pooling image features $x_{\text{siglip}}$ through an attention pooling head in SigLIP (Zhai et al., 2023). We randomly select 1,500 images from the MSCOCO test set and visualize these representations using t-SNE in Figure 3 (b). Notably, the representations from VILA-U are relatively distant from SigLIP, while TokLIP shares nearly the same distribution as SigLIP2. This aligns with TokLIP’s stronger retrieval performance demonstrated in Section 4.2. These analyses further show that our pipeline learns more similar semantics to pre-trained CLIP ViTs, thus achieving superior comprehension performance.
@@
Pretrained on large-scale image-text pairs, CLIP (Radford et al., 2021; Tschannen et al., 2025; Lin et al., 2024a; Gao et al., 2024; Ma et al., 2025b) has achieved remarkable success in capturing high-level semantic features of images, making it highly effective for feature extraction in multimodal large language models (LLMs). Numerous studies (Liu et al., 2023; Zhu et al., 2023; Bai et al., 2023; Lu et al., 2024; Lin et al., 2024b; Zhou et al., 2024b, 2025) leverage CLIP’s continuous features by combining them with text tokens in LLMs and training on vision-language data to enhance performance in comprehension tasks. However, while CLIP excels at encoding high-level semantics, it often overlooks low-level structures and textural details, which are crucial for visual generation tasks. Models such as SEED-X (Ge et al., 2024), DreamLLM (Dong et al., 2023), and MetaMorph (Tong et al., 2024) incorporate CLIP’s continuous features with text tokens for autoregressive prediction but still rely on diffusion models (Podell et al., 2023) for generation inference to compensate for the lack of low-level details. In contrast, TokLIP proposes a unified tokenizer that integrates both low-level features and high-level semantics, enabling direct tokenization with CLIP encoders and allowing for more effective handling of visual generation tasks with VQGAN decoder.
@@ 4 Experiments
[…]
EVA02-B/16 (Fang et al., 2024) Merged-2B 224 74.7 58.74 42.15
CLIP-L/14 (Radford et al., 2021) WIT400M 224 75.5 56.34 36.51
MetaCLIP-L/14 (Xu et al., 2023) MetaCLIP400M 224 76.2 60.00 43.81
VILA-U (Wu et al., 2024b) COYO700M 256 73.3 62.56 45.37
QLIP (Zhao et al., 2025) DataComp1B 256 74.3 55.62 38.91
TokLIP-B Mix 125M 256 76.4 64.06 48.46
OpenCLIP-ViT-L/14 (Cherti et al., 2023) LAION2B 336 75.3 63.36 46.51
SigLIP-B/16 (Zhai et al., 2023) WEBIL10B 384 76.5 67.74 49.90
SigLIP-So/14 (Zhai et al., 2023) WEBIL10B 384 83.2 70.20 52.00
VILA-U (Wu et al., 2024b) COYO700M 384 78.0 - -
QLIP (Zhao et al., 2025) DataComp1B 384 79.1 60.86 43.00
TokLIP-L Mix 80M 384 80.0 68.00 52.87
SigLIP-B/16 (Zhai et al., 2023) WEBIL10B 512 79.1 68.72 50.55
SigLIP2-B/16 (Tschannen et al., 2025) WEBIL10B 512 81.2 71.20 55.20
SigLIP2-L/16 (Tschannen et al., 2025) WEBIL10B 512 83.5 72.10 55.20
TokLIP-XL Mix 70M 512 80.8 69.40 53.77